Actions
RunPlan20150619 » History » Revision 3
« Previous |
Revision 3/4
(diff)
| Next »
Drielsma, François, 19 June 2015 16:57
June 19th, 2015 (Evening Run)¶
Aim: Tracker alignment data
Daily schedule¶
June | 19th | |
Time | 10pm - 6am | |
MOM | V. Blackmore | |
Lead shifter | C. Heidt | |
Shifter 2 | P. Francini | |
BLOC | P. Hanlet | |
SOC | C. Heidt | |
Experimenter | E. Overton, M. Uchida | |
Physics Shifter | C. Rogers | |
Beam? | Yes | |
Additional Personnel |
Shift Monitoring¶
Shifters should keep a careful eye on:- Water chiller temperature outside the control room (bottom of the red LEDs). If this reaches 25C, immediately power down the magnets and inform the MOM.
- Magnet water temperatures.
- Log any abnormal changes in the eLog
- If yellow temperature alarm sounds, immediately power down the magnets and inform the MOM.
- Magnet water flows.
- Log any abnormal changes in the eLog
- If yellow alarm sounds, immediately power down the magnets and inform the MOM.
- EMR stopping the DAQ arming
- Power cycle the VME crate by following the instructions here
Shift plan:¶
- 8:45pm
- Shifters + BLOC arrive
- Hall and beam line to be closed and prepared for beam
- 10pm
- Check magnet chilled water temperatures in archiver
- Take 500 spills of Configuration 1 and calculate ratio of TOF1 to TOF2 triggers (i.e. attempt to quantify transmission)
- Take 500 spills of Configuration 2 and calculate ratio of TOF1 to TOF2 triggers (i.e. attempt to quantify transmission)
- Continue data taking of beam setting with largest transmission, minimum 50k accepted triggers at TOF2
- If transmission of beams is similar and 50k triggers at TOF2 is easily achieved within shift, take 50k triggers at TOF2 of the alternative configuration. Otherwise, prefer higher momentum (configuration 1)
Data Requested¶
- (3, 300) muon beam
- Trigger from TOF1
- 50k triggers at TOF2
Beamline Configurations¶
- We expect to run with 2V beamloss and the beam bump active.
- DS is off
- Beam settings via Run Control
Configuration 1¶
Setting for 300 MeV/c however upstream is set to the max available current (a bit lower than typically coming from Marco's table):
Number of pulses | Beamloss | Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) | RunControl tag |
# | V | MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # | ||
Pi+/Mu+ | 512 | 508.75 | 333 | 82 | 109.21 | 199.75 | 122.30 | 437.83 | - | 124.71 | 221.83 | 297.48 | 197.36 | 202.43 | 306.46 | 261.93 | 300+No_DS |
Configuration 2¶
Setting for 270 MeV/however upstream is set to the max available current, which than corresponds to about 270 MeV/c into the Cooling Channel:
Number of pulses | Beamloss | Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) | RunControl tag |
# | V | MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # | ||
Pi+/Mu+ | 512 | 508.75 | 299.26 | 82 | 109.21 | 199.75 | 122.30 | 437.83 | - | 112.49 | 199.23 | 267.19 | 177.23 | 180.04 | 272.52 | 232.88 | 270+No_DS |
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Updated by Drielsma, François over 8 years ago · 3 revisions